# Holographic dictionary for Lifshitz and hyperscaling violating black holes

### v1

@@ Appendix A Mass formula, scaling relations and extended first law
×$\quad\times\bigg{(}1+\frac{(d-2)^{2}(d-\theta+z-2)(d-\theta+z-1)k}{16\pi GP(d-\theta+z-3)^{2}}\left(\frac{r_{0}}{r_{F}}\right)^{\frac{2\theta}{d-1}}\left(\frac{4GS}{\omega_{k,d-1}r_{F}^{\theta}}\right)^{-\frac{2}{d-\theta-1}}$
@@
$M(S,Q,P,r_{F},r_{0}^{-\gamma})\,$ $=\alpha^{-\theta}M(\alpha^{\theta}S,\alpha^{\theta-\theta/(d-1)}Q,\alpha^{-2\theta/(d-1)}P,\alpha r_{F},r_{0}^{-\gamma})\quad$ (38)
@@
×$\quad\times\bigg{(}r_{h}^{-2(d+z-1)}(d-\theta+z-3)^{2}\big{(}(z+1)r_{h}^{2(d+z)}+q^{2}(z-1)r_{h}^{2\theta+4}\big{)}+(d-2)^{2}kL^{2}(z-1)\bigg{)}\,,$
@@
The quantity $U$ vanishes if $\theta=0$, and $U_{2}$ is zero if both $\theta=0$ and $z=1$, in which case $\Theta$ reduces to $-r^{d}\omega_{k,d-1}/d$.
@@
Since the pressure depends via $\Lambda_{0}$ on $L$, $r_{F}$ and $r_{0}$, due to (12), the extended first law can also be written in terms of variations of these length scales
@@ Appendix C Euler equation from a generalized Cardy formula
Since the Cardy formula expresses the entropy as a function $\tilde{S}=\tilde{S}(\tilde{E},V,C),$ with central charge $C$ defined in (59) and the “volume” of the circle given by $V=2\pi R$, we can explicitly compute the temperature, pressure and chemical potential associated with the central charge using the definitions that follow from the first law (20)

### v2

@@ Appendix A Mass formula, scaling relations and extended first law
$\quad\times\bigg{(}1+\frac{(d-2)^{2}(d-\theta+z-2)(d-\theta+z-1)k}{16\pi GP(d-\theta+z-3)^{2}}\left(\frac{r_{0}}{r_{F}}\right)^{\frac{2\theta}{d-1}}\left(\frac{4GS}{\omega_{k,d-1}r_{F}^{\theta}}\right)^{-\frac{2}{d-\theta-1}}$
@@
$M(S,Q,P,r_{F},r_{0}^{-\gamma})\,$ $=\alpha^{-\theta}M(\alpha^{\theta}S,\alpha^{\theta-\theta/(d-1)}Q,\alpha^{-2\theta/(d-1)}P,\alpha r_{F},r_{0}^{-\gamma})\quad\$(38)
@@
$\quad\times\bigg{(}r_{h}^{-2(d+z-1)}(d-\theta+z-3)^{2}\big{(}(z+1)r_{h}^{2(d+z)}+q^{2}(z-1)r_{h}^{2\theta+4}\big{)}+(d-2)^{2}kL^{2}(z-1)\bigg{)}\,$,
@@
The quantity $U$ vanishes if $\theta=0$, and $U_{2}$ is zero if both $\theta=0$ and $z=1$, in which case $\Theta$ reduces to $r^{d}\omega_{k,d-1}/d$.
@@
Since the pressure depends via $\Lambda_{0}$ on $L$, $r_{F}$ and $r_{0}$, due to (12) in the main text, the extended first law can also be written in terms of variations of these length scales
@@ Appendix C Euler equation from a generalized Cardy formula
Since the Cardy formula expresses the entropy as a function $\tilde{S}=\tilde{S}(\tilde{E},V,C),$ with central charge $C$ defined in (59) and the “volume” of the circle given by $V=2\pi R$, we can explicitly compute the temperature, pressure and chemical potential associated with the central charge using the definitions that follow from the first law